### v1

# Beyond Specialization: Assessing the Capabilities of MLLMs in Age and Gender Estimation

Abstract
Multimodal Large Language Models (MLLMs) have recently gained immense popularity. Powerful commercial models like ChatGPT-4V and Gemini, as well as open-source ones such as LLaVA, are essentially general-purpose models and are applied to solve a wide variety of tasks, including those in computer vision. These neural networks possess such strong general knowledge and reasoning abilities that they have proven capable of working even on tasks for which they were not specifically trained. We compared the capabilities of the most powerful MLLMs to date: ShareGPT4V, ChatGPT, LLaVA-Next in a specialized task of age and gender estimation with our state-of-the-art specialized model, MiVOLO. We also updated MiVOLO and provide details and new metrics in this article. This comparison has yielded some interesting results and insights about the strengths and weaknesses of the participating models. Furthermore, we attempted various ways to fine-tune the ShareGPT4V model for this specific task, aiming to achieve state-of-the-art results in this particular challenge. Although such a model would not be practical in production, as it is incredibly expensive compared to a specialized model like MiVOLO, it could be very useful in some tasks, like data annotation.
1 Introduction
Figure 1: An example of evaluated models predictions. The image illustrates output of specialized model (MiVOLO) and different MLLMs. MiVOLO makes predictions based on the face and body crops. Other models make predictions based on prompt and image of body crop.
The rapid development of large multimodal language models (MLLMs or LMMs) has been noteworthy, particularly those integrating language and vision modalities (LVMs). Their advancement is attributed to their high accuracy, generalization capability, reasoning skills, and robust performance with out-of-distribution data. These versatile models excel not only as AI assistants but also in handling unforeseen tasks beyond their initial training scope. The impact of MLLMs is profound, evolving so swiftly that it raises questions about the relevance of specialized models in certain areas. Moreover, there is an increasing interest in using MLLMs for specific computer vision tasks, such as object segmentation, and incorporating them into complex pipelines, such as instruction-based image editing.
In July 2023, we introduced the state-of-the-art model for gender and age determination, MiVOLO [13]. While developing its second version, we explored how competitive MLLMs are in this domain. Initially, ChatGPT4V [28] could not process images of individuals, but this capability was enabled during our project, prompting us to conduct preliminary tests. The results were highly encouraging, leading to a comprehensive evaluation of these neural networks’ potential, including leading open-source solutions like LLaVA [24] [22] and ShareGPT4V [5], which is also based on LLaVA.
We pursued several goals in these experiments:
  • We aimed to compare the best general-purpose MLLMs with specialized models and understand their capability to replace them. Despite the huge difference in computational costs and speed, for some tasks, this is not crucial. This includes tasks such as labeling new data or filtering old datasets.
  • We were interested in what results could be achieved if an MLLM was fine-tuned for a specific task on a large target dataset. With the same motivations, many tasks require maximum accuracy and do not require fast inference.
  • Since the nature of specialized models and large general-purpose models is fundamentally different, it was reasonable to expect that such experiments could shed more light on the strengths and weaknesses of both approaches.
For the experiments, we tried to measure all SOTA MLLM models: LLaVA 1.5 and LLaVA-NeXT, ShareGPT4V, and ChatGPT4V. We were unable to measure the newly released Gemini Ultra, as it outright refused to work with images of people.
Since these experiments were conducted as part of improving MiVOLO and comparisons are mostly made with it, we also briefly present the main changes of the new version in this article.
Figure 1 demonstrates an example of work of evaluated models and figure 2 provides a graphical representation.
Figure 2: Graphical representation of the study focused on comparison specialized models for age and gender estimation and multimodal models.
2 Related Works
Age and gender estimation models. Different researchers approach the problem of recognizing a person’s age in various ways and typically address it using classical machine learning methods, CNNs or transformer-based models, primarily relying on face crops as input data.
Authors of [37, 47, 2] employ classical machine learning methods to tackle the regression problem. [3] utilizes ResNet34 to determine age through ranking of results of multiple binary classification models.
The work [16] approaches the recognition task of facial attributes, such as age and gender, as multiclass and binary classification problems, respectively, employing CNN to generate predictions. [34] also deals with classification problems, adapting MobileNet to simultaneously predict age and gender. [35] trains CNN to classify gender and age as part of multiple facial analysis tasks using multi-task learning. [25] utilizes two GoogLeNet models to predict age: an age classifier and an age regressor. [46] represents age as a convex combination of two other numbers and employs a CNN to predict the weights for these numbers.
The publication [39] leverages CNN and GCN [12] to extract semantic features alongside structural information from the face image for age prediction.
Neural networks based on transformers with an attention mechanism [41], commonly employed for various natural language processing (NLP) tasks, are also widely utilized for CV tasks. Authors of [42, 11] utilize not only CNNs, but also attention mechanisms, directing the model to focus on features relevant for age estimation. [42] selects the most informative age-specific patches for age estimation. [11] uses transformer to aggregate the sequence of embeddings extracted by CNN and further utilyze the aggregated feature vector for age estimation. MiVOLO [13] employs a transformer to estimate age and gender using face and body crops as input data. In this paper, we introduce MiVOLOv2, which outperforms all specialized models mentioned above.
Model Train Dataset Test Dataset Age MAPE, % $\downarrow$ Age MAE $\downarrow$ Age CS@5, % $\uparrow$ Gender Acc, % $\uparrow$
FP-Age [20] IMDB-clean IMDB-clean - 4.68 63.78 -
MiVOLOv1 LAGENDA IMDB-clean 11.43 [19.25] 4.09 [6.66] 69.72 [47.53] 99.55 [96.74]
224x224 [13] train LAGENDA 13.19 [28.88] 3.99 [7.41] 71.27 [49.64] 97.36 [93.57]
MiVOLOv2 LAGENDA IMDB-clean 11.01 [17.40] 3.97 [6.03] 71.16 [52.11] 99.68 [97.56]
384x384 train LAGENDA 12.06 [24.64] 3.65 [6.16] 74.48 [55.90] 97.99 [95.69]
Table 1: Comparison of MiVOLOv1 and MiVOLOv2 Performance. Both models are evaluated in two modes: $face+body$ and $body\ only$. The results for the $body\ only$ mode are shown in brackets.
Multimodal Models. Pre-trained vision-language models like CLIP [32] are extensively utilized in computer vision tasks. They notably improve performance across various downstream tasks by effectively matching text and images [33, 10, 26].
Some works [43, 19, 38] use the CLIP pre-trained model for age estimation through visual and text embeddings matching. However, vision-language models still encounter difficulties in understanding instructions, capturing context, and adapting to unseen tasks. Consequently, many researchers are investigating ways to transfer the capabilities of more powerful LLMs into the visual domain, leading to the development of multimodal large language models (MLLM) [1, 18, 45, 24, 48, 31, 44]. Other researchers are leveraging the robust abilities of MLLM for multimodal understanding and generation to address vision tasks [15, 6, 17, 27, 9].
In this paper, we aim to evaluate the capabilities of MLLMs in age and gender estimation tasks.
Examining the potential of multimodal ChatGPT (ChatGPT4V[29]), authors of [7] assess its aptitude in predicting various facial attributes and executing face recognition tasks. With zero training the model outperformed a specialized model in age recognition, but performed less effectively in gender classification. We also compare the ChatGPT4V[29] model with a specialized trained model and also consider open source state-of-the-art MLLMs such as ShareGPT4V[5], LLaVa-1.5[24], LLaVa-NeXT[23]. We also fine-tune a MLLM for the task of gender and age estimation to compare it with a specialized model.
3 MiVOLOv2
In this section, we briefly describe the changes in the new version of MiVOLO that we use in subsequent experiments to compare with MLLM models as an anchor specialized model.
3.1 Evaluation metrics
We utilize the same evaluation metrics as those referenced in previous study [13]:
  • Mean Absolute Error (MAE) for age estimation.
  • Cumulative Score at 5 (CS@5) for age estimation.
  • Accuracy for gender prediction.
Additionally, we have slightly modified the Mean Average Percentage Error (MAPE) in this study as follows:
$\mbox{MAPE}=\frac{1}{n}\sum_{i=1}^{n}\left|\frac{y^{pred}_{i}-y^{gt}_{i}}{y^{gt}_{i}+\epsilon}\right|$ (1)
We chose $\epsilon=1$ in the denominator for two main reasons: to prevent division by zero and to mitigate excessively high percentage errors in cases involving infants. For instance, employing $\epsilon=0.083$ (approximately one month) would result in disproportionately large errors for infants, thereby significantly biasing the MAPE.
3.2 Datasets
We use the same data as before, additionally extending the train dataset by about 40%, so it now contains more than 807,694 samples: 390,730 male and 416,964 female. To do so, we primarily used our production pipelines and some open-source data, like LAION-5B [36]. We selected mostly images where MiVOLOv1 made significant mistakes. We also balanced the right tail of the distribution as much as possible – previously, the training part still suffered from imbalance after about 70 years, despite our efforts to compensate for this.
3.3 Experiment Details
The training procedure did not change significantly. We conducted many experiments with additional training stage image augmentations but retained only one new augmentation – face blurring, to imitate social network filters or effects from smartphone cameras. In the first stage of training, where we train a single-input model that uses only faces, we apply this blur only slightly, with a 5% random chance. In the second stage, with double-input, we apply blurring with significantly increased parameters, up to almost complete blurring of the face, and with a 70% probability. We also introduce a model trained with new input size 384x384 instead of previously used 224x224. This shows much better results for the in-the-wild domain. For this new model, we decrease dropout and drop-path rates to only 0.1, because the dataset is now large and diverse enough, so less noise caused by dropout works better.
3.4 Results
We briefly present the results in Table 1, which show significant improvements. The results for MiVOLOv2 establish new state-of-the-art results for specialized models. For comparisons with MLLM models, see the following section. Table 2 provides a comparison of results using the Cross-Age Celebrity Dataset (CACD2000) [4] test split.
Model Age MAE $\downarrow$ Age CS@5, % $\uparrow$
ResNet-50 [30] 3.96 -
MiVOLOv1 224 [13] 4.09 70.73
MiVOLOv2 384 3.89 73.26
Table 2: Comparison of models using the CACD test split. Both versions of MiVOLO are evaluated in face + body mode.
4 MLLM Models vs. Specialized Model
4.1 Benchmarks
A multitude of benchmarks is available for age or gender estimation. In this study, we concentrated on those offering full-body images, when possible, and containing both age and gender labels.
Hence, we selected two datasets:
  • IMDB-clean dataset [21] used in the MiVOLO article [13]. The dataset was enhanced with body bounding boxes associated with facial pairs. It comprises 183,886 training images, 45,971 validation images, and 56,086 test images. We use only the test images for our evaluation.
  • LAGENDA [13], a dataset well-balanced in terms of age and gender attributes, features face-body pairs with ground truth obtained through human annotations via weighted voting. It includes proprietary training and validation parts, and an open-source test part, containing 67,159 images from the Open Images Dataset [14] featuring 84,192 individuals aged from 0 to 95.
These datasets are of exceptionally high quality and exhibit significant diversity.
Initially, the OpenAI API imposed a limit of 100 requests per day for the gpt-4-vision-preview models, which has since been increased to 1,500 requests per day. Due to this limitation, we randomly selected a small subset from LAGENDA to run with ChatGPT4V and included it in our comparison. We selected 200 random samples for each age group at intervals of 5 years (e.g., 0-5, 5-10, etc.) for ages $[0;90]$. However, $>21\%$ of these images had to be removed because ChatGPT refused to provide answers. As a result, this dataset contains 3,062 samples. We refer to this dataset as NanoLAGENDA.
We opted for LAGENDA over IMDB to minimize the risk that MLLMs would provide correct answers not through age and gender estimation but because of its familiarity with famous individuals, well-known movies, etc. On the other hand, LAGENDA, annotated by human annotators, does not have actual ground truths for labels. Nevertheless, we chose it because the risk of this drawback is lower. We will demonstrate later in this article that our assumption of LAGENDA’s labels being very close to actual ground truths on average, based on measurements from [13], is valid.
Additionally, to slightly offset the drawback of annotated ages, we compiled a very small dataset of 104 samples from social networks like Instagram, primarily with real ground truth answers (we know the actual ages) for ages $[0;105]$ and very challenging samples, which typically result in large human prediction errors. This dataset is entirely manual and is intended solely to double-check our conclusions. We refer to this small set as the Wild104 dataset. We cannot publish it because we do not own the photos.
We also used the Adience [8] benchmark to compare specialized MiVOLO models with the multimodal approaches mentioned in section 2 and our fine-tuned models 4.3. The dataset consists of 26,580 facial images. Annotations include age labels from eight age group classes and labels for the binary gender classification task.
4.2 Method
Base Model Body or Face Crop Age MAE $\downarrow$ Age CS@5, % $\uparrow$ Gender Acc, % $\uparrow$
LLaVa-v1.5-7b Body 4.52 69.84 99.06
ShareGPT4V-7b 0.4 epoch Body 3.87 75.93 99.45
ShareGPT4V-7b 1 epoch Body 3.94 75.34 99.53
ShareGPT4V-13b Body 3.93 75.42 99.54
ShareGPT4V-7b Face 4.32 72.98 97.56
Table 3: Comparison of fine-tuned MLLMs on LAGENDA benchmark.
We used the same prompt for all models. However, the full version was necessary only for ChatGPT; for the sake of an honest comparison, we applied the same for open-source models as well. In our tests, this did not influence the answers.
Prompt for MLLMs
A typical answer looks like: [female; 40]
We set the temperature for all models to 0.0. For ChatGPT, we additionally set the parameter ”seed” to 1234 and ”n” to 1. The latter is necessary due to reports that just the seed and temperature are not sufficient to ensure deterministic and reproducible results for vision model via API.
However, for different models, a zero temperature setting can lead to various issues due to the nature of LLMs. ChatGPT4V might occasionally fail to provide an answer, typically for queries that are either difficult or do not pass the safety system checks. We decided to remove such samples, although it gives ChatGPT4V a slight advantage. LLaVA [24], LLaVA-NeXT [22], and LLaVA-based ShareGPT4V [5] may sometimes return an age range instead of a specific age. Since we have sampling disabled and the temperature is set to 0.0, we cannot workaround this; thus, we take the midpoint of the range in such cases. However, the number of such samples is very low, and this does not significantly impact the results.
For all models, we attempted to use the maximum possible resolution with the goal to measure the maximum possible performance without taking into accound speed of inference. Thus:
  • For ChatGPT, we used full-sized original crops to allow the model to split them into tiles (if large enough), which gives ChatGPT another slight advantage.
  • For LLaVA 1.5 and ShareGPT4V, we used the original 336x336 resolution.
  • For LLaVA-NeXT, we utilized the new Dynamic High Resolution technique, although its maximum resolution is still restricted to 672x448 or 448x672.
Model Input Age Acc, % $\uparrow$ MAE, % $\downarrow$ Gender Acc, % $\uparrow$
OridinalCLIP [19] Face 61.2 ± 4.2 0.47 ± 0.6 -
L2RCLIP [43] Face 66.2 ± 4.4 0.36 ± 0.5 -
MiVOLOv1 224x224 [13] Body & face 68.68 0.345 96.5
MiVOLOv2 384x384 Body & face 69.43 0.333 97.39
ShareGPT4V 7B fine-tuned Body 66.7 0.349 95.65
ShareGPT4V 7B fine-tuned with faces Face 67.95 0.338 96.63
Table 4: Comparison of models using Adience benchmark. MAE here is calculated on classification labels.
Model Input Age MAPE, % $\downarrow$ Age MAE $\downarrow$ Age CS@5, % $\uparrow$ Gender Acc, % $\uparrow$
LLaVA 1.5 7B [24] Entire body 16.86 7.59 $\pm$ 7.22 48.79 99.38
Body w/o face 43.18 18.44 $\pm$ 17.38 25.71 94.85
LLaVA-NeXT 34B[23] Entire body 13.73 6.20 $\pm$ 5.78 55.40 99.51
Body w/o face 23.03 9.58 $\pm$ 8.51 39.67 97.82
ShareGPT4V 7B [5] Entire body 16.71 7.16 $\pm$ 6.48 53.24 99.44
Body w/o face 25.80 11.16 $\pm$ 10.45 39.07 97.24
ChatGPT4V [29] Entire body 12.12 4.66 $\pm$ 4.51 68.10 98.43
Body w/o face 22.56 7.82 $\pm$ 7.52 49.15 93.02
MiVOLOv2 384x384 Body & face 11.61 4.33 $\pm$ 4.47 69.90 97.71
Body w/o face 21.82 7.19 $\pm$ 7.67 49.14 95.61
ShareGPT4V 7B fine-tuned Entire body 10.95 4.22 $\pm$ 3.99 72.09 99.51
Body w/o face 20.53 7.64 $\pm$ 7.35 49.95 97.48
ShareGPT4V 13B fine-tuned Entire body 11.30 4.26 $\pm$ 4.18 73.34 99.44
Body w/o face 19.90 7.40 $\pm$ 7.72 54.01 98.10
Table 5: Comparison of performance on NanoLAGENDA benchmark. Models are evaluated with different type of input information. Bold indicates the best model performance running with all available information about the person. Underline shows the best performance running without faces.
4.3 LLaVA finetune
In this section, we explore the fine-tuning of a general-purpose multimodal network, namely LLaVA, for age and gender recognition tasks.
Building on insights from the MiVOLO paper, simultaneous training for both tasks has shown to be advantageous. Consequently, our training approach mirrors the evaluation methodology described in the preceding section.
Initially, the only available version for research was LLaVa-v1.5’s training code in open source, guiding our choice of starting point. Data conversion utilized the same prompt as for ChatGPT, employing a single crop as the input image. Various experiments using both whole body and face crops were conducted, with whole body crops yielding marginally superior results for gender and age recognition tasks.
Further, ShareGPT4V, which is derived from LLaVa code and demonstrates slightly improved metrics, became the source of checkpoints from the Hugging Face hub, used with minor modifications to the LLaVa code. The ShareGPT4V training code was not accessible at the time.
Another experimental direction focused on training for direct gender classification and age prediction, rather than text prediction, using linear layers along with MSE and cross-entropy losses, paralleling the approach used for LLaMa sequence classification tasks [40]. This approach is not mentioned in the results due to bad performance.
The optimal hyperparameters identified through our experiments are as follows: learning rate = $2e^{-6}$, LoRA disabled (full fine-tuning), per_device_train_batch_size = 32, num_train_epochs = 1, checkpoint every 450 iterations, warmup_ratio = 0.03, lr_scheduler_type ”cosine”
Table 3 provides a comparative overview of the fine-tuned MLLMs. Note that the best-performing model in Table 3 (ShareGPT4V-7b 0.4 epoch) is referred to as ShareGPT4V 7B fine-tuned in subsequent sections, representing the optimal checkpoint achieved.
The finest results were obtained using the ShareGPT4V-7b model trained with whole body crops. Notably, the best metrics were observed at the 900th iteration checkpoint, approximately 40% through one epoch, suggesting an early stop with low learning rate strategy might be beneficial and shows that MLLMs, as generalist models, are easy to fine-tune with smaller data amount.
Important to mention, that after a few iterations, the training loss stabilizes at around 0.32, and further training steps may lead to overfitting. This is corroborated by the observation that later iterations yield slightly diminished results. Additionally, extended training may impair the model’s assistant capabilities, restricting responses to the trained format. This limitation could potentially be mitigated by diversifying the training data beyond age estimation tasks. It is also worth noting that testing on the Lagenda test set requires approximately 2.5 hours on 8 A6000 NVidia GPUs, a significant duration relative to the training time of about 11 hours for one epoch. While other iteration intervals may yield superior results, our study focused on evaluating every 450th iteration to optimize training time and costs.
Future developments could explore dual-crop inputs (body + face as separate images), as seen in the original MiVOLO model. However, the feasibility of training existing models with multiple images per input remains an open question.
4.4 Results
The table 4 presents a comparison of specialized MiVOLO models and multimodal approaches using the Adience benchmark. Following the methodology of [13], we mapped regression predictions to the nearest intervals (classes).
Table 5 displays analogous results for a randomly sampled subset of NanoLAGENDA, including evaluations of ChatGPT4V.
Figure 3 illustrates the relationship between MAE and age for NanoLAGENDA, with age intervals set at 5-year steps. Interestingly, ChatGPT’s performance closely parallels that of models trained specifically on our dataset, particularly underperforming in the 25 to 55 age range—a notably common age group in the dataset.
Table 6 reports outcomes for the full-sized LAGENDA and IMDB datasets for applicable models.
Table 7 shows results for the Wild104 benchmark, reaffirming prior findings with real-world ground truth labels and highlighting a different visual domain.
Figure 3: Relationship between MAE and age group across different models tested on the NanoLAGENDA benchmark.
Model Test Dataset Age MAPE, % $\downarrow$ Age MAE $\downarrow$ Age CS@5, % $\uparrow$ Gender Acc, % $\uparrow$
MiVOLOv2 IMDB-clean 11.01 3.97 71.16 99.68
384x384 LAGENDA 12.06 3.65 74.48 97.99
LLaVA-NeXT 34B IMDB-clean 16.04 5.66 59.77 99.15
vanilla [23] LAGENDA 16.97 5.19 62.17 99.47
ShareGPT4V 7B IMDB-clean 12.07 4.40 70.28 99.47
fine-tuned LAGENDA 11.47 3.52 79.66 99.44
Table 6: Comparison of performance on IMDB and LAGENDA benchmarks.
Model Age MAPE, % $\downarrow$ Age MAE $\downarrow$ Age CS@5, % $\uparrow$ Gender Acc, % $\uparrow$
ChatGPT4V [29] 19.95 7.07 $\pm$ 6.05 48.08 91.35
LLaVA-NeXT 34B vanilla [23] 22.16 9.23 $\pm$ 8.45 42.31 96.15
MiVOLOv2 384x384 19.82 6.26 $\pm$ 6.17 53.67 96.17
ShareGPT4V 7B fine-tuned 19.36 7.01 $\pm$ 6.67 53.85 95.19
ShareGPT4V 13B fine-tuned 18.27 6.79 $\pm$ 6.42 57.69 97.12
Table 7: Comparison of performance on the Wild104 benchmark: observed standard deviations indicate models instability in average errors.
Figure 4: Failure case. Through manual analysis of ChatGPT’s gender misclassification, we have discovered that it sometimes errs with examples where men have long hair. This is a jokingly artificial illustrative example of such failure cases. Both MiVOLO and the fine-tuned as well as vanilla LLaVA models remain stable.
5 Conclusions
This study aimed to assess the efficacy of cutting-edge specialized models in comparison to Multimodal Large Language Models (MLLMs) for age and gender estimation tasks.
Our findings reveal a nuanced view. MLLMs, despite not being explicitly trained for facial or bodily analyses to deduce personal attributes, exhibit exceptional capabilities. Notably, models such as ChatGPT stand out by harnessing vast amounts of visual information and training on extensive datasets, demonstrating significant proficiency in tasks beyond their original design. Our research indicates that with minor adjustments, MLLMs can achieve or even surpass the performance of specialized models, suggesting a potential shift towards versatile, general-purpose networks in computer vision. This hints at a future where Multimodal Small Language Models (MSLMs) could potentially replace current specialized models. The flexibility of language models, reducing the need for task-specific models, offers significant advantages for a wide range of applications, especially in scenarios where computational resources and inference speed are not primary concerns. However, in specific data subsets, a well-designed specialized model may still outperform a fine-tuned MLLM. It’s important to note, for the time being, the computational cost of MLLMs cannot be directly compared to that of specialized models — the difference can span thousands of times. Our analysis identified ChatGPT4V as the most precise MLLM for age estimation across numerous benchmarks, despite encountering challenges such as occasional refusal to process images with significant losses ($>21\%$ for our data) and the need for Dynamic High Resolution, which demands a higher computational budget. While this comparison may not be entirely equitable, it highlights the model’s capabilities in ’maximum power mode’. Among open-source alternatives, LLaVA-NeXT 34B leads in this area. The same time, specialized model MiVOLO surpasses all general-purpose MLLMs in age estimation. However, for certain data segments and metrics, fine-tuned specialized versions of LLaVA prove more effective. Such fine-tuned MLLMs present a promising solution for many tasks where computational cost is not a primary concern. Compared to the intensive and expert-driven training required for MiVOLO, fine-tuning an MLLM is considerably simpler, requiring only the same dataset as the specialized model and minimal expertise. Original hyperparameters and losses can be used.
The study highlights the superior performance of MLLMs in gender identification tasks even without any fine-tuning, surpassing that of specialized models. This emphasizes the significance of high-level feature recognition and contextual understanding in this task, where nearly all MLLMs excel. However, ChatGPT-4V stands out for its subpar performance, possibly due to an overemphasis on certain features like hair, which might be influenced by its training data or safety mechanisms. For visualizations, refer to Figure 4. The opaque nature of its development process hampers definitive conclusions. While MLLMs face challenges in maintaining task-specific performance consistency, experiencing notable fluctuations with minor image variations, specialized models show greater stability.
In conclusion, despite current challenges such as high inference costs and other obstacles that may impede their deployment in production environments, the future of technology appears distinctly multimodal. Advances in neural network research and hardware development are progressively making the practical application of multimodal language models more feasible, enhancing their abilities while reducing operational costs. It is expected that the use of versatile and powerful MLMs in production settings will grow, steadily reducing the dependence on specialized models for a wide range of tasks.